OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

y=-10 , x= 6

OpenStudy (anonymous):

u can use subsitution method

OpenStudy (anonymous):

how did you get that?

OpenStudy (uri):

@mr.singh Stop giving direct answers,It's against Coc.

OpenStudy (anonymous):

x-y=16 .. x=16+y... put in second equation nd solve for y

OpenStudy (uri):

x - y = 16 2x + y = 2 We add the both equation remember -y +y will be cancelled...Do it and show me what you get @jaderbrown

OpenStudy (anonymous):

I got 6

OpenStudy (anonymous):

3x=18 3 3 x=6

OpenStudy (uri):

x - y = 16 2x + y = 2 ======== 3x=18 x=18/3 x=6 Correct

OpenStudy (anonymous):

now what?

OpenStudy (uri):

Now take the original 2nd eq and put the value of x 2x + y = 2 So,2(6)+y=2 Right?

OpenStudy (anonymous):

yes

OpenStudy (uri):

Hm so, 2(6)+y=2 12+y=2 Shift the 12 to right side sign changes So y=2-12 y=-10

OpenStudy (anonymous):

thank you!

OpenStudy (uri):

You're welcome :)